OpenStudy (anonymous):

Solveeeeeeeeeeeeeee if p(x) = px^7 + qx^3 + rx - 5 where p , q , r are constants , if p(-7) = 7 then p(7) = ?

OpenStudy (anonymous):

not enough info to give a real number for f(7)

OpenStudy (anonymous):

however we can give the answer in terms of p,q,r

OpenStudy (anonymous):

how? and it's p(7) not f(7)

OpenStudy (perl):

this looks like an odd function

OpenStudy (anonymous):

yeah p(7)

OpenStudy (anonymous):

lets start with plugging in p\(-7)=7

OpenStudy (anonymous):

I don't know if this works.. p(x) = px^7 + qx^3 + rx - 5 p(-7) = p(-7)^7 + q(-7)^3 + r(-7) - 5 = 7 p(-7)^7 + q(-7)^3 + r(-7) = 12 - [ p(7)^7 + (q(7)^3 + r(7)] = 12 p(7)^7 + (q(7)^3 + r(7) = -12 So, p(7) = p(7)^7 + q(7)^3 + r(7) - 5 Plug in the results for p(7)^7 + (q(7)^3 + r(7) , I think we can get the answer..

OpenStudy (anonymous):

\[-7^7p-7^3q-7r-5=7 \implies -7^7p-7^3q-7r=12\] \[p(7)=7^7p+7^3q+7+5=-12+5=7\]

OpenStudy (anonymous):

@RolyPoly Yes, i m done same procedure, but i at end i m stuck

OpenStudy (anonymous):

some how we got the answer because of symmetry

OpenStudy (anonymous):

@msingh Stuck? where?

OpenStudy (anonymous):

wait u have done some where wrong @RolyPoly

OpenStudy (anonymous):

p(7)=7

OpenStudy (anonymous):

p(x) = px^7 + qx^3 + rx - 5 p(-7) = p(-7)^7 + q(-7)^3 + r(-7) - 5 = 7 p(-7)^7 + q(-7)^3 + r(-7) = 12 - [ p(7)^7 + (q(7)^3 + r(7)] = 12 p(7)^7 + (q(7)^3 + r(7) = -12 How can u get minus sign common, it's impossible

OpenStudy (perl):

@Jonask theres one mistake in your calculations

OpenStudy (perl):

p(7)=7^7p+7^3q+7-5=-12-5=-17

OpenStudy (perl):

the function is neither even nor odd ,

OpenStudy (anonymous):

oh yeah i made a mistake

OpenStudy (anonymous):

@Jonask I don't think giving out answer makes you cool p(-7)^7 + q(-7)^3 + r(-7) = 12 For the first term: p (-7)^3 = p(-7)(-7)(-7)...(-7) = -p(7)^7 For the second term:q(-7)^3 = q (-7)(-7)(-7) = -q(7)^3 So, take out -1 as the common factor. - [ p(7)^7 + (q(7)^3 + r(7)] = 12

OpenStudy (perl):

no worries, i wont take away your medal :)

OpenStudy (anonymous):

p(x) = px^7 + qx^3 + rx - 5 p(-7) = p(-7)^7 + q(-7)^3 + r(-7) - 5 = 7 p(-7)^7 + q(-7)^3 -7r = 12

OpenStudy (anonymous):

@Jonask Giving out answer doesn't make you cool

OpenStudy (perl):

@RolyPoly thats rude rolypoly, he gave an explanation

OpenStudy (anonymous):

thanks for the knowledge ,i know also its part of the openstudy code of conduct 4 give me

OpenStudy (perl):

whats openstudy code of conduct 4,

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (perl):

I believe Jonask guided the person

OpenStudy (anonymous):

i have only one medal otherwise i wud have given to all of u one one, its not something to fight on, everyone know the answer some answer it some not.

OpenStudy (anonymous):

I'm not here for medals.

OpenStudy (perl):

do you guys agree the answer is p(7)= -17

OpenStudy (anonymous):

yes

OpenStudy (perl):

by the way, even if p(7) = 7, that is not enough to make the function symmetric. because it could be a coincidence, you need all elements in your domain p(-x) = p(x) for even p(-x) = - p(x) for odd. now do you see why its neither even nor odd. its a sum of an odd function plus an even function

OpenStudy (anonymous):

@perl how it came -17

OpenStudy (perl):

one sec

OpenStudy (perl):

p(x) = px^7 + qx^3 + rx - 5 p(-7) = p(-7)^7 + q(-7)^3 + r(-7) - 5 = 7 p(-7)^7 + q(-7)^3 + r(-7) = 12 - [ p(7)^7 + (q(7)^3 + r(7)] = 12 so we know that p(7)^7 + (q(7)^3 + r(7) = -12 now find p(7) p(7) = p(7)^7 + (q(7)^3 + r(7) -5 , and substitute from above = ( -12) - 5

OpenStudy (anonymous):

thank u